OpenStudy (mony01):

Determine whether the series is absolutely convergent, conditionally convergent, or divergent.

OpenStudy (mony01):

\[\sum_{n=1}^{\infty}(-1)^{n}\frac{ \arctan n }{ n^9 }\]

OpenStudy (anonymous):

is it decreasing? do the terms go to zero?

OpenStudy (anonymous):

if that is not clear, \[\lim_{n\to\infty}\arctan(n)=\frac{\pi}{2}\]

OpenStudy (anonymous):

and therefore it converges for sure, at least conditionally, as the series alternates and the terms go to zero that is all you have to check for an alternating series